Evaluate


∫ √tanx. sec4x dx

π

4



Also integrate the definite integral using MATLAB command.


1
Expert's answer
2021-12-13T16:09:06-0500

"I=\int^{\pi /4}_0 \sqrt{tanx}sec^4x dx=\int^{\pi /4}_0 \sqrt{tanx}(1+tan^2x)sec^2xdx"


"tanx=t,sec^2x=dt"


"I=\int(t^{1/2}+t^{5/2})dt=2t^{3/2}/3+2t^{7/2}/7=2tan^{3/2}x/3+2tan^{7/2}x/7|^{\pi/4}_0="


"=2\/3+2\/7=20\/21"


MATLAB code

syms x
expr=sqrt(tanx)*(secx)^4;
F=int(expr,0,pi/4)